OpenStudy (anonymous):

Find a positive inverse for 3 modulo 40. That is find a positive integer s such that 3s=1(mod 40)

terenzreignz (terenzreignz):

27 works, I guess

OpenStudy (anonymous):

haha that is teh right answer but how did u get that?

terenzreignz (terenzreignz):

I don't know, guesswork, which I guess, is not the right way to go about it. Perhaps you would enlighten me?

OpenStudy (anonymous):

u need to use linear combination. but can u explain me ur logic?

terenzreignz (terenzreignz):

yeah, I just used trial and error... 40.... 40+1 divisible by 3? no 80.....80+1 divisible by 3? yes, it must be 81/3 then :)

terenzreignz (terenzreignz):

And I don't know what linear combination is :P

OpenStudy (anonymous):

hahah ok thanks :)

OpenStudy (experimentx):

programmatically i found 27 too

OpenStudy (anonymous):

ya its the answer but like i dont know how to solve ut. like they have this weird method

OpenStudy (experimentx):

3*s - 1 must be greater than 40 which means s>14

OpenStudy (experimentx):

3s-1 must have zero at end therefore s must be 17, 27, 37, 47, ..

OpenStudy (anonymous):

okkkk i see u just use logic

OpenStudy (experimentx):

i used computer to find it first ..

OpenStudy (experimentx):

on the other hand, we have 3s - 1 should be divisible by 4 too

OpenStudy (experimentx):

or, 3s = 1mod(4) 3s = 1mod(10)

OpenStudy (experimentx):

from, 3s = 1mod(4) we have s = 3,7,11,15,19,23,27,31, ... hence, the first intersection occurs at s = 27 so this is the answer

OpenStudy (anonymous):

ohhh i seee i like this method

OpenStudy (experimentx):

well, i'm still not pretty much convinced if you can use this method ... algebra should be more fundamental than this ... i guess

OpenStudy (anonymous):

THANKS :DDDDDDDDD

OpenStudy (experimentx):

yw .. if i find any better, i'll update

OpenStudy (anonymous):

you want to solve \(3x\equiv 1( \text{mod }40)\) one way to do this is use the euclidean algorithm

OpenStudy (experimentx):

i'm interested .. please go on

OpenStudy (anonymous):

ya that is method i need to use

OpenStudy (anonymous):

in this case you are lucky, because it only takes one step

OpenStudy (anonymous):

you get \(40=13\times 3+1\)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

this tell you \(1=1\times 40-13\times 3\) and so the inverse of 3 is -13

OpenStudy (anonymous):

which of course is the same mod 40 as 27

OpenStudy (anonymous):

didnt follow the last step

OpenStudy (anonymous):

you are lucky in this case that you get it right away. if you wanted compute the inverse of 7 mod 40 it might take more steps was the answer clear?

OpenStudy (anonymous):

ok i thought maybe not

OpenStudy (anonymous):

liek abt the inverse

OpenStudy (anonymous):

you want \(3\times x\equiv 1(40)\) which by definition means 40 divides \(3x-1\)

OpenStudy (anonymous):

which in turn means \(3x-1=40k\) for some \(k\) or \(1=40j+3x\)

OpenStudy (anonymous):

ok got that part

OpenStudy (anonymous):

\(j\) and \(x\) are called the "bezout coeffients" so that is what you are looking for

OpenStudy (anonymous):

by the euclidean algorithm, divide 40 by 3 as needed until you get a remainder of 1. in this case it only takes one step, so this one is very easy. so we have \[40=13\times 3+1\] solving for 1 gives \[1=40-13\times 3\] so the coefficient for 3 is -13

OpenStudy (anonymous):

that means the inverse of 3 is -13 but generally since you are working mod 40 your universe is positive integers between 0 and 39, so you would say -13 = 40-13 = 27 and that is your inverse

OpenStudy (anonymous):

ohhhhh i get it. Like my book never explained what inverse was

OpenStudy (anonymous):

lol what book?

OpenStudy (anonymous):

text book

OpenStudy (anonymous):

YAY that was awesome SAT :D

OpenStudy (experimentx):

well ... thanks for new method ...

OpenStudy (anonymous):

we can also do this the donkey way, which might be easier and make more sense

OpenStudy (anonymous):

start with \(3x\equiv 1(40)\) and you want to divide by 3, but you cannot so keep adding 40 until you can

OpenStudy (anonymous):

hahah i never heard of teh donkey way lol

OpenStudy (anonymous):

\(3x\equiv 1(40)\) \[3x\equiv 41(40)\] still can't divide by 3 \[3x\equiv 81(40)\] ahh now we can divide! \[3x\equiv 81(40)\] \[x\equiv 27(40)\]

OpenStudy (anonymous):

this way is in fact rather easy, especially if you have small numbers and can get it right away

OpenStudy (anonymous):

in fact if i was going to do this on an exam, this is probably the method i would use

OpenStudy (anonymous):

OK thanks sattelite. Sorry os is freezing my comp

OpenStudy (anonymous):

yw, hope it is clear